Are these vectors linearly independent? • [3 4 0 2],[4 − 2 6 14],[0 8 − 6 − 11]
Answer #1

Answers let a[3402] + b[4-2 6 14] rc{08-6-11]=10000 3a +464, 42-2b, 6b-6c, 2a +146-11c] = [0000] 3a74b=0H9-26:0, 6b-60=0,29 +